1- Find the mesure of an angle if its measure is 3 degrees less than twice the measure of its supplement

Answer by 303795(602)

If the angle is x then its supplement is 180-x.
so x+3=2(180-x)
x+3=360-2x
3x=357
x=119
supplement = 61